# SRA-CL: Semantic Retrieval Augmented Contrastive Learning for Sequential Recommendation

### v4

@@ B.3 Implementation Details
All experiments are conducted with a single V100 GPU. The embedding size for all methods is set to 64 for a fair comparison. We use a training batch size of 256 and employ the Adam optimizer with a learning rate of 0.001. The dropout rate is set to 0.5 for both the embedding layer and the hidden layers across all datasets. Following previous studies [33], we set the maximum sequence length to 20 for all datasets. The early stopping is applied if the metrics on the validation set do not improve over 10 consecutive epochs. Our method is model-agnostic and can be applied to any sequential recommendation model. The transformer backbone mentioned in Sec. 2.2 comprises two layers, each with two attention heads. For the LLM, we select DeepSeek-V3, a robust large language model that demonstrates exceptional performance on both standard benchmarks and open-ended generation evaluations. For detailed information about DeepSeek, please refer to their official website [^3] https://github.com/deepseek-ai/DeepSeek-V3. Specifically, we utilize DeepSeek-V3 by invoking its API [^4] https://api-docs.deepseek.com/. To reduce text randomness of the LLM, we set the temperature $\tau$ to 0 and the top-$p$ to 0.001. For the text embedding model $\mathcal{M}$, we use the pre-trained SimCSE-RoBERTa [^5] https://huggingface.co/princeton-nlp/sup-simcse-roberta-large from Hugging Face. Identical settings are used for baselines that involve LLMs and text embeddings to ensure fairness.

### v5

@@ B.3 Implementation Details
All experiments are conducted with a single V100 GPU. The embedding size for all methods is set to 64 for a fair comparison. We use a training batch size of 256 and employ the Adam optimizer with a learning rate of 0.001. The dropout rate is set to 0.5 for both the embedding layer and the hidden layers across all datasets. Following previous studies [yang2024sequential], we set the maximum sequence length to 20 for all datasets. The early stopping is applied if the metrics on the validation set do not improve over 10 consecutive epochs. Our method is model-agnostic and can be applied to any sequential recommendation model. The transformer backbone mentioned in Sec. 2.2 comprises two layers, each with two attention heads. For the LLM, we select DeepSeek-V3, a robust large language model that demonstrates exceptional performance on both standard benchmarks and open-ended generation evaluations. For detailed information about DeepSeek, please refer to their official website [^3] https://github.com/deepseek-ai/DeepSeek-V3. Specifically, we utilize DeepSeek-V3 by invoking its API [^4] https://api-docs.deepseek.com/. To reduce text randomness of the LLM, we set the temperature $\tau$ to 0 and the top-$p$ to 0.001. For the text embedding model $\mathcal{M}$, we use the pre-trained SimCSE-RoBERTa [^5] https://huggingface.co/princeton-nlp/sup-simcse-roberta-large from Hugging Face. Identical settings are used for baselines that involve LLMs and text embeddings to ensure fairness.